write the equation y^2- 8y - 8x = 8 in standard form. Fill in the blanks and sketch the graph.
A)vertex:
B)Directrix:
C)AOS:
D)p=
E)LR=
F)focus:

Answer by lynnlo(4176)

(1)(y-4)^2=8(x+3)
(A)(-3,4)
(B)x=-5
(C)Y=4
(D)P=
(E)lr=
(F)(-1,4)